CAT Chat Minutes

April 10, 1998

Issues that require action and/or follow-up

For the long-term schedule, we are looking at providing 5000 user bean hours during FY '99, with 2 long shutdowns, one planned for mid-December to mid-January and another in May. For the intervention times, we are looking at a two-day studies/maintenance period every other week. The week between, we would take 1 shift for studies. The two-day periods would give us the opportunity to complete some maintenance projects, if needed, and give us more time for some studies. We will also be able to use the 48 hours for PSS revalidations, which we will have to do every 6 months. The cycles will begin on Wednesdays, not Mondays or Fridays.

User Comments:

*BESSRC - It is our experience that, when handling large volumes of users, the one-day (Monday)for studies is very favorable for set up. When experiments run a week, there is one day for set up in between.

Having the studies time during the week is better than coming in on the weekend to help get users set up.

DND - With the situation we experienced, it would be nice to get the repair work done and be able to run again in a couple of days as opposed to being down for the entire run.

*The last time we had work done on our PSS, the two days was really a problem for SRI CAT. Is something being done to shorten the time it takes to complete this ?

Yes, we are working on simplifying the validation and thereby shortening the time. At the present time, with software safety systems, when the code has to be rewritten, the entire system must be revalidated.

*How many days will be set aside for top-off mode?

We will work with you. The ASD and XFD staffs have done some studies that we will get together and go over, and then, with your participation, we will plan further studies on this mode. We will give you further information before we begin this.

*Is top-off like trickle?

The intent is to maintain the current in the SR to within 0.1 to 1% of maximum current. Later on, with experience, it may be within 0.01%.

*Is there any information about the new parking in the 435 area?

We have submitted a request for a new parking area here; we will check on the progress.

*Last week when running my spectroscopy, I had great data. This week I tried twice to run this experiment, and had poor quality. Has anyone else seen this?

Yes, Sector 13 has experienced similar problems also, the BPM had no trace. SRI had to chase down the beam.

M. Ramanathan will look into this. Please let us know when you see these things.

*When will the card-keys be installed on the truck lock doors?

We have issued metal keys to all the CATs for these doors, but we had no plans of installing card-keys on these doors.

*We have noticed a significant delay in the pickup of chemical waste from DND LOM. On occasion they have not been successful in getting into the areas that have card key access. What can be done?

While delays are not unusual, we are concerned about this problem and will address it.

*Is there a log of previously posted messages from the control room somewhere? It would be nice for users to be able to access any archived messages.

We will try to provide this.

Information/Follow-up Items

APS Operational Status Update

The machine has been running well. Earlier this week we had problems with power supplies and rf. This run ends at midnight on Sunday 4-12-98. We have a five-week shutdown. User operations begins on May 18, 1998. Cycle 1998-3 is on the web. Currently we have seven Tuesdays for shielding verification. We will use two. The others will be used for top-off studies. We will define the tests for these runs with user participation. Monday all the ratchet doors will be open. We will require entrance to each of them at some point. Work is planned for some FEs.

Information

There is a power shutdown scheduled for one week from tomorrow. It will only affect 431.

During the shutdown, we will be installing four remote shutter controllers.

You will see a lot of activity next week at 5 ID, please do not interfere. We will keep you informed.

The ring road lane painting from parking lots and stop signs at the 431 truck lock area will begin the week of April 13, 1998.

The cash machine now gives $5.00 bills.

May 2, 1998, there will be a site-wide power reduction. We will be limited to 4 megawatts or less. You may be asked to shut things down. Maintenance will be performed on the main power lines feeding the laboratory.

During the next couple of weeks, helios and saline will go down for upgrades. The following week oxygen and ozone will also be upgraded. ECT has asked to run the security software. We ask that you let your FC know when would be the best time to have ECT run this software on your machines.